OpenStudy (anonymous):

[\sqrt[5]{16}\times 2^{\frac{1}{5}}\]

OpenStudy (tommynaut):

Is that the fifth root of 16 multiplied by 2 to the power of one fifth? You could convert the fifth root of 16 into the fractional index form to help simplify things.

OpenStudy (phi):

\[\sqrt[5]{16}\times 2^{\frac{1}{5}}\] you can write that as \[ 16^\frac{1}{5} \cdot 2^{\frac{1}{5}}\] and write 16 as 2^4 \[ \left(2^4\right)^\frac{1}{5} \cdot 2^{\frac{1}{5}}\]